OpenStudy (daniellelovee):

A password contains exactly 5 letters. How many passwords are possible if letters cannot be used more than once?

OpenStudy (daniellelovee):

@Directrix

OpenStudy (anonymous):

65

Directrix (directrix):

Why 65? How did you get it?

OpenStudy (anonymous):

5*26=130 130/2 = 65

OpenStudy (anonymous):

oops

OpenStudy (daniellelovee):

I got 7,893,600

OpenStudy (anonymous):

we have \(26\) letters. for the first letter, there are \(26\) possibilities; for the second, having used one letter, there are only \(25\); for the third, then, using similar reasoning, there are \(24\); and so on. the number of possible passwords is then: $$26\cdot25\cdot24\cdot23\cdot22=7893600$$

OpenStudy (anonymous):

5= password max 26= numbers of letters in alphabet 130= possibilities including re-occuring numbers /2 = No re-occurances

Directrix (directrix):

A password contains exactly 5 letters. Are the 5 letters of the password distinct? And, are the possible passwords all of length 5 letters? @Daniellelovee

OpenStudy (anonymous):

this is also written using factorials as: $$\frac{26\cdot25\cdot24\cdots1}{21\cdots1}=\frac{26!}{(26-5)!}=\ _{26}P_5$$

OpenStudy (anonymous):

Correct! @oldrin.bataku

OpenStudy (daniellelovee):

no but oldrin just told me that my answer was correct thank you @Directrix

Directrix (directrix):

Nowhere is it said that the passwords have to be of length 5 letters. letters cannot be used more than once? --> Does not mean the letters are used once.

OpenStudy (daniellelovee):

mhhh true

Directrix (directrix):

>A password contains exactly 5 letters. Okay, I assumed that the password was a specified one. This is a generic any 5 letter password adhering to the conditons. Yet, do we know that the passwords to be formed from the letters are 5-letter passwords?

OpenStudy (daniellelovee):

is not specified

OpenStudy (daniellelovee):

is my response correct? @Directrix